OpenStudy (anonymous):

How far downstream would a swimmer land if the river is 423m and the swimmer swims at a rate 2m/s and then current is perpendicular to the swimmer is 1.50 m/s?

OpenStudy (jamesj):

The river is 423m wide?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

Ok, at a rate of 2 m/s, now many seconds does it take to cross the river?

OpenStudy (anonymous):

174 sec

OpenStudy (jamesj):

No. speed = distance/time, v = d/t. Hence t = d/v If d = 423 m and v = 2 m/s, then t = d/v = 423/2 = 211.5 seconds.

OpenStudy (jamesj):

Now, in that time, how far down stream does the current carry the swimmer?

OpenStudy (anonymous):

yes, i have no idea about this, my teacher didn't teach us this part!

OpenStudy (jamesj):

Well, the swimmer is carried downstream at what speed?

OpenStudy (anonymous):

1.50 m/s

OpenStudy (jamesj):

Yes and how far will that be in t = 211.5 seconds?

OpenStudy (anonymous):

exactly

OpenStudy (jamesj):

speed = distance/time. Hence distance = speed x time